### v1

@@ 27 Dims 54 Dims 81 Dims 108 Dims 135 Dims 162 Dims 189 Dims 216 Dims 243 Dims
Mental Disability vs Non-Disabled 1,43 0,97 0,45 0,53 0,90 0,31 1,11 0,14 1,28 -0,06 1,37 -0,22 1,21 -0,42 1,00 -0,58 0,84 -0,78 0,64
Mental Disability vs Physical Disability 0,92 1,45 -0,53 1,47 -0,55 1,50 -0,58 1,52 -0,60 1,50 -0,58 1,49 -0,57 1,46 -0,54 1,46 -0,54 1,46 -0,55
Non-Disabled vs Physical Disability -1,06 1,00 0,05 1,39 -0,33 1,50 -0,44 1,57 -0,51 1,60 -0,55 1,64 -0,58 1,67 -0,61 1,70 -0,64 1,74 -0,68
American vs Arab -0,97 0,31 0,67 0,86 0,11 1,09 -0,12 1,26 -0,29 1,32 -0,35 1,38 -0,41 1,46 -0,49 1,62 -0,65 1,81 -0,83
American vs Chinese -0,56 -0,53 0,03 -0,42 0,15 -0,41 0,15 -0,32 0,24 -0,27 0,29 -0,14 0,43 0,05 0,52 0,46 0,11 1,22 -0,65
American vs Mexican -1,07 0,40 0,67 0,91 0,16 1,15 -0,08 1,28 -0,22 1,36 -0,29 1,41 -0,34 1,50 -0,44 1,65 -0,59 1,82 -0,76
Arab vs Chinese 0,53 -0,77 -0,25 -0,87 -0,34 -1,30 -0,77 -1,37 -0,85 -1,40 -0,88 -1,41 -0,88 -1,41 -0,89 -1,47 -0,94 -1,57 -1,04
Arab vs Mexican -0,13 0,10 0,03 0,15 -0,02 0,11 0,02 0,03 0,10 0,09 0,04 0,11 0,02 0,14 -0,01 0,15 -0,02 0,15 -0,01
Chinese vs Mexican -0,65 0,84 -0,18 1,14 -0,49 1,34 -0,69 1,39 -0,74 1,44 -0,78 1,44 -0,79 1,46 -0,81 1,52 -0,86 1,61 -0,95
Buddhist vs Christian 0,80 0,96 -0,16 1,03 -0,23 1,01 -0,21 1,05 -0,25 1,16 -0,36 1,28 -0,48 1,30 -0,50 1,46 -0,66 1,65 -0,85
Buddhist vs Hindu 0,00 0,39 -0,38 0,45 -0,44 0,46 -0,45 0,44 -0,43 0,48 -0,48 0,60 -0,59 0,55 -0,55 0,58 -0,58 1,00 -1,00
Buddhist vs Jewish -1,66 -0,95 0,71 -0,36 1,29 -0,21 1,44 0,00 1,65 0,28 1,38 0,52 1,13 0,62 1,03 1,03 0,62 1,53 0,13
Buddhist vs Muslim -1,60 -0,85 0,75 -0,37 1,23 -0,24 1,35 -0,20 1,40 0,06 1,54 0,42 1,18 0,56 1,03 0,97 0,63 1,47 0,12
Christian vs Hindu -0,73 -0,60 0,13 -0,65 0,08 -0,60 0,12 -0,67 0,06 -0,80 -0,07 -0,83 -0,11 -0,91 -0,18 -1,09 -0,37 -1,17 -0,44
Christian vs Jewish -1,71 -1,39 0,32 -1,27 0,44 -1,09 0,62 -1,00 0,71 -0,98 0,73 -0,95 0,76 -0,91 0,80 -0,84 0,87 -0,55 1,17
Christian vs Muslim -1,67 -1,33 0,34 -1,20 0,47 -1,07 0,60 -1,07 0,61 -1,05 0,62 -0,94 0,73 -0,88 0,79 -0,81 0,86 -0,53 1,14
Hindu vs Jewish -1,58 -1,07 0,50 -0,85 0,73 -0,60 0,98 -0,41 1,17 -0,22 1,36 -0,11 1,46 0,04 1,54 0,45 1,13 0,84 0,73
Hindu vs Muslim -1,53 -1,00 0,53 -0,77 0,77 -0,61 0,93 -0,55 0,99 -0,37 1,16 -0,16 1,37 0,02 1,51 0,42 1,12 0,80 0,74
Jewish vs Muslim -0,18 0,01 0,17 0,04 0,14 -0,05 0,13 -0,19 0,00 -0,18 0,00 -0,06 0,12 -0,02 0,17 0,00 0,18 -0,01 0,17
Heterosexual vs LGBT -1,33 -1,34 0,00 -1,35 -0,02 -1,37 -0,04 -1,41 -0,08 -1,48 -0,15 -1,56 -0,22 -1,74 -0,41 -1,74 -0,40 -1,72 -0,38
Average gain 3,86 4,03 4,08 4,24 4,00 3,45 2,98 0,10 -3,30

### v2

@@ 27 Dims 54 Dims 81 Dims 108 Dims 135 Dims 162 Dims 189 Dims 216 Dims 243 Dims
Mental Disability vs Non-Disabled 1.43 0.97 0.45 0.53 0.90 0.31 1.11 0.14 1.28 -0.06 1.37 -0.22 1.21 -0.42 1.00 -0.58 0.84 -0.78 0.64
Mental Disability vs Physical Disability 0.92 1.45 -0.53 1.47 -0.55 1.50 -0.58 1.52 -0.60 1.50 -0.58 1.49 -0.57 1.46 -0.54 1.46 -0.54 1.46 -0.55
Non-Disabled vs Physical Disability -1.06 1.00 0.05 1.39 -0.33 1.50 -0.44 1.57 -0.51 1.60 -0.55 1.64 -0.58 1.67 -0.61 1.70 -0.64 1.74 -0.68
American vs Arab -0.97 0.31 0.67 0.86 0.11 1.09 -0.12 1.26 -0.29 1.32 -0.35 1.38 -0.41 1.46 -0.49 1.62 -0.65 1.81 -0.83
American vs Chinese -0.56 -0.53 0.03 -0.42 0.15 -0.41 0.15 -0.32 0.24 -0.27 0.29 -0.14 0.43 0.05 0.52 0.46 0.11 1.22 -0.65
American vs Mexican -1.07 0.40 0.67 0.91 0.16 1.15 -0.08 1.28 -0.22 1.36 -0.29 1.41 -0.34 1.50 -0.44 1.65 -0.59 1.82 -0.76
Arab vs Chinese 0.53 -0.77 -0.25 -0.87 -0.34 -1.30 -0.77 -1.37 -0.85 -1.40 -0.88 -1.41 -0.88 -1.41 -0.89 -1.47 -0.94 -1.57 -1.04
Arab vs Mexican -0.13 0.10 0.03 0.15 -0.02 0.11 0.02 0.03 0.10 0.09 0.04 0.11 0.02 0.14 -0.01 0.15 -0.02 0.15 -0.01
Chinese vs Mexican -0.65 0.84 -0.18 1.14 -0.49 1.34 -0.69 1.39 -0.74 1.44 -0.78 1.44 -0.79 1.46 -0.81 1.52 -0.86 1.61 -0.95
Buddhist vs Christian 0.80 0.96 -0.16 1.03 -0.23 1.01 -0.21 1.05 -0.25 1.16 -0.36 1.28 -0.48 1.30 -0.50 1.46 -0.66 1.65 -0.85
Buddhist vs Hindu 0.00 0.39 -0.38 0.45 -0.44 0.46 -0.45 0.44 -0.43 0.48 -0.48 0.60 -0.59 0.55 -0.55 0.58 -0.58 1.00 -1.00
Buddhist vs Jewish -1.66 -0.95 0.71 -0.36 1.29 -0.21 1.44 0.00 1.65 0.28 1.38 0.52 1.13 0.62 1.03 1.03 0.62 1.53 0.13
Buddhist vs Muslim -1.60 -0.85 0.75 -0.37 1.23 -0.24 1.35 -0.20 1.40 0.06 1.54 0.42 1.18 0.56 1.03 0.97 0.63 1.47 0.12
Christian vs Hindu -0.73 -0.60 0.13 -0.65 0.08 -0.60 0.12 -0.67 0.06 -0.80 -0.07 -0.83 -0.11 -0.91 -0.18 -1.09 -0.37 -1.17 -0.44
Christian vs Jewish -1.71 -1.39 0.32 -1.27 0.44 -1.09 0.62 -1.00 0.71 -0.98 0.73 -0.95 0.76 -0.91 0.80 -0.84 0.87 -0.55 1.17
Christian vs Muslim -1.67 -1.33 0.34 -1.20 0.47 -1.07 0.60 -1.07 0.61 -1.05 0.62 -0.94 0.73 -0.88 0.79 -0.81 0.86 -0.53 1.14
Hindu vs Jewish -1.58 -1.07 0.50 -0.85 0.73 -0.60 0.98 -0.41 1.17 -0.22 1.36 -0.11 1.46 0.04 1.54 0.45 1.13 0.84 0.73
Hindu vs Muslim -1.53 -1.00 0.53 -0.77 0.77 -0.61 0.93 -0.55 0.99 -0.37 1.16 -0.16 1.37 0.02 1.51 0.42 1.12 0.80 0.74
Jewish vs Muslim -0.18 0.01 0.17 0.04 0.14 -0.05 0.13 -0.19 0.00 -0.18 0.00 -0.06 0.12 -0.02 0.17 0.00 0.18 -0.01 0.17
Heterosexual vs LGBT -1.33 -1.34 0.00 -1.35 -0.02 -1.37 -0.04 -1.41 -0.08 -1.48 -0.15 -1.56 -0.22 -1.74 -0.41 -1.74 -0.40 -1.72 -0.38
Average gain 3.86 4.03 4.08 4.24 4.00 3.45 2.98 0.10 -3.30